OpenStudy (anonymous):

in a group of 120 students, 40 students like listening to rock and country music. If 63 students like rock and 47 students like country music, how many do not like listening to either one?

OpenStudy (anonymous):

50 students

OpenStudy (sepeario):

40+63+47= 150-120=30 students. Well atleast that's how I do it...

OpenStudy (anonymous):

what is this process?